# Scalable Multi-Agent Path Finding using Collision-Aware Dynamic Alert Mask and a Hybrid Execution Strategy

### v1

@@ 3.1 Approach Summary
where $v^{k}_{0}=s_{k}$, $v^{k}_{\tau_{k}}=g_{k}$, and $\pi_{\theta}$ is a parameterized RL policy trained to minimize path length and local collision risk over a planning horizon $H$. No information about other agents is exchanged during this stage. Let $\tau_{i}$ denote the set of makespan of each agent and $\tau=\{\tau_{1},\dots,\tau_{n}\}$ be the set of all such makespans.
@@
S4: Decentralized Replanning
Upon receiving a collision alert, agent $a_{c_{k}}$ truncates its path at $v^{c_{k}}_{t_{j-r}}$, denoted as
@@
The agent then generates a new path segment $\rho_{c_{k}}^{\prime}$ from $v^{c_{k}}_{t_{j-r}}$ to its goal $g_{c_{k}}$ by applying its RL policy $\pi_{\theta}$. This replanning incorporates new constraints derived from the alert. For static obstacle avoidance (S4.1), the policy is typically derived by incorporating a fixed constraint, wherein the agent is supposed to avoid forbidden cells, as captured in Eq. 6, where $\Delta_{c}$ represents the set of static cells identified in the alert, and a segment of the new path is generated as:
$\rho_{c_{k}}^{\prime}=\pi_{\theta}(v^{c_{k}}_{t_{j-r}},g_{c_{k}}\mid v^{c_{k}}_{t_{i}}\notin\Delta_{c},i\geq j-r).$ (6)
@@
Alternatively, for dynamic obstacle avoidance (S4.2), the replanning must account for the predicted movements of other agents involved in the collision. In this case, the RL policy is conditioned on the sub-paths of the conflicting agents
$\rho_{c_{k}}^{\prime}=\pi_{\theta}(v^{c_{k}}_{t_{j-r}},g_{c_{k}}\mid v^{c_{k}}_{t}\neq v^{c_{l}}_{t},v^{c_{l}}_{t}\in\rho_{c_{l}},$
$\forall l\neq k,a_{c_{l}}\in A_{c},\forall t\in[t_{j-r},t_{j+r}]).$ (7)
@@
The information guiding the replanning–whether it constitutes minimal details such as static obstacle constraints (Eq. 6), or more detailed sub-path information about colliding agents’ paths treated as dynamic obstacles (Eq. 7)–is integrated into the RL agent’s decision-making process (by modifying its state representation, more details in the Section 3.2) to guide it towards a conflict-free solution. The agent’s new complete trajectory $\rho^{new}_{c_{k}}$ is formed by concatenating the initial segment with the newly planned one, given by
$\rho^{new}_{c_{k}}=\rho_{c_{k}|t_{j-r}}\|\rho_{c_{k}}^{\prime}.$ (8)
This updated trajectory is submitted to the central controller. This iterative cycle of detection (S2), control (S3), and selective replanning (S4) continues until a globally conflict-free solution is achieved ($C(\rho,\tau)=\emptyset$).
Our hybrid MAPF framework manages a precise flow of information, essential to support our claims of reduced information sharing and adaptive planning. The process begins with fully decentralized agent planning (S1), where each RL-driven agent uses only its local information before submitting its intended path $\rho_{k}$ to a central coordinator. This coordinator, leveraging its global observation of all submitted paths, performs collision detection (S2) to identify the set of all potential conflicts $C(\rho,\tau)$. Following detection, the central control module (S3) issues targeted alerts; it selects a specific agent $a_{c_{k}}$ and determines the rewind point $t_{j-r}$ from which the agent must replan, effectively issuing an alert $\mathcal{A}(c)$. S3 also governs the level of information to be shared for the subsequent decentralized replanning (S4). Initially, the alerted agent $a_{c_{k}}$ attempts to resolve the conflict using minimal information, typically by treating its own conflicting path segment $\Delta_{c}$ as a static obstacle (S4.1). If this localized, low-information approach proves insufficient, S3 then shares additional information with the agent–specifically, a subset of the path of other directly conflicting agents–for a more informed replan (S4.2). This tiered strategy—from purely local information in S1, to targeted alerts from S3, and then to progressively detailed yet still localized conflict information for S4—ensures that inter-agent information sharing is demand-driven and sparse. Agents communicate their revised plans $\rho^{new}_{c_{k}}$ back to the coordinator, and this iterative cycle of detection and selective, adaptive replanning continues until a globally conflict-free solution is achieved ($C(\rho,\tau)=\emptyset$). This methodology substantiates our framework’s ability to ensure feasibility while operating with significantly reduced information exchange.
@@ 3.2 Policy Representation
In our proposed framework, we address the challenges of decentralized multi-agent pathfinding in dynamic environments by integrating collision awareness directly into the agent’s observation. Our approach leverages an RL-based decentralized planner that receives a multi-channel grid observation.
Observation Space: Each agent’s observation is a tensor $\mathbf{s}\in\mathbb{R}^{H\times W\times 4}$, where $H$ and $W$ are the height and width of the grid. The four channels correspond to: (a) ObstacleMap - A binary map (1 if a static obstacle is present, 0 otherwise). (b) AgentMap - A binary map marking the agent’s current position. (c) GoalMap - A binary map indicating the goal position. (d) AlertMask - A dynamic collision alert map (initially all zeros, updated during execution). This alert mask is updated online by a centralized collision detection module that simulates timely collision alerts. In addition, the observation includes low-dimensional features comprising a unit vector pointing from the agent’s current position to the goal and the Euclidean distance.
Action Space: The agent operates in a discrete action space, $\mathcal{A}=\{0,1,2,3,4\}$, corresponding to movements in the four cardinal directions (e.g., Up, Down, Left, Right, WAIT).
Reward Structure Our reward function, similar to PRIMAL (Sartoretti et al. 2019), promotes efficient, collision-free navigation and discourages unproductive behaviors. Agents receive +20 for reaching the goal. Penalties include: -3 for collisions with obstacles, -2 for timeout, -0.02 per time step taken (to encourage efficiency), and -0.1 for a ‘WAIT’ action, a penalty that encourages path progression over stationary waiting. An additional -0.05 penalty is applied if the agent is near a dynamic obstacle.
@@
The collision detection module (S2) functions as a deterministic, rule-based system. It takes the set of all current agent trajectories $\rho=\{\rho_{1},\dots,\rho_{n}\}$ and their makespans $\tau$ as input. For each timestep $t=\{0,1,...,T_{M}\}$ upto the maximum makespan ($T_{M}$), the module systematically scans for two primary types of conflicts: vertex conflicts (where $v^{k}_{t}=v^{\ell}_{t}$ for $k\neq\ell$) and edge conflicts (where $(v^{k}_{t},v^{k}_{t+1})=(v^{\ell}_{t+1},v^{\ell}_{t})$ for $k\neq\ell$). Upon detecting a conflict, characterized as $c=(t,v,A_{c})$ where $A_{c}$ is the set of agents involved, S2 reports this conflict $c$ to the S3 control module for resolution. The detection process for all conflicts is computationally efficient, with a time complexity of $O\bigl(\sum_{k}|\rho_{k}|\bigr)$ per cycle, linear in the sum of all agent path lengths.
@@ 4 Experimental Setup
In this section we first describe how we train the per‐agent navigation policy used in stage S1, then detail the maps, problem instances, and competing planners used to evaluate stages S2–S4, and define the performance metrics used to evaluate.
@@
We train a parametrized RL model $Q_{\theta}(s,a)$ using the DDQN algorithm(Van Hasselt, Guez, and Silver 2016), incorporating prioritized experience replay (PER) and $\varepsilon$‑greedy exploration, to enable an agent to navigate to a specified goal in the presence of both static and dynamic obstacles. Each dynamic obstacle possesses a hidden goal and follows precomputed trajectories, executing only valid moves; this allows for the simulation of online collision alerts during inference from the S3 stage. Training is performed for 30 00030\,000 episodes on an $11\times 11$ maze grid, with each episode capped at a maximum of $T_{\max}=50$ steps. The curriculum for environmental complexity is scheduled as follows: during the first 500 episodes, static‑obstacle density $\rho_{s}=0.10$ and dynamic‑obstacle count $n_{d}=0$; from episode 500 to 2999, $\rho_{s}=0.10$ and $n_{d}=1$; from episode 3000 to 5999, $\rho_{s}=0.20$ and $n_{d}=2$; and for episode 6000 onward, $\rho_{s}=0.30$ and $n_{d}=4$.
Input observations are normalized to zero mean and unit variance before being fed to the network. Transitions $(s_{t},a_{t},r_{t},s_{t+1})$ are stored in a prioritized experience replay buffer of size $10^{6}$. At each learning step, a mini‑batch of $128$ transitions is sampled. The agent selects actions using an $\varepsilon$‑greedy policy, where $\varepsilon$ decays from $\varepsilon_{0}=1.0$ to $0.01$ over the course of training according to $\varepsilon_{t+1}=\max(0.01,0.999\,\varepsilon_{t})$. Crucially, during both exploration and exploitation (e.g., when calculating the target Q-value), only actions $a_{t}$ deemed valid by an action mask $m_{t}\in\{0,1\}^{|\mathcal{A}|}$ are considered for efficient exploration (Damani et al. 2021b).
The Q-network parameters $\theta$ are updated via gradient descent using the Adam optimizer with a learning rate $\alpha=3\times 10^{-4}$. The discount factor is $\gamma=0.97$. A separate target network, with parameters $\theta^{-}$, provides stable targets and is updated by copying $\theta$ every 300 steps. Updates minimize the prioritized, importance‑weighted Bellman error. The target value $y_{t}$ and TD-error $\delta_{t}$ are computed as: $y_{t}=r_{t}+\gamma\,Q_{\theta^{-}}\bigl(s_{t+1},\arg\max_{a^{\prime}}Q_{\theta}(s_{t+1},a^{\prime})\text{ s.t. }m_{t}(a^{\prime})=1\bigr),$ $\delta_{t}=y_{t}-Q_{\theta}(s_{t},a_{t}),$ The loss function is then calculated as $L(\theta)\;=\;\mathbb{E}_{i\sim p(i)}\bigl[w_{i}\,\delta_{i}^{2}\bigr],$ where the sampling probability $p(i)\propto|\delta_{i}|^{\alpha}$ ($\alpha$ is a hyperparameter for PER) and $w_{i}$ are importance‑sampling weights that correct for the bias introduced by prioritized sampling.
In addition to the DDQN model, a Proximal Policy Optimization (PPO) based model was also trained; details regarding its architecture (including the one used for DDQN model) and training procedure, along with comparative performance, are provided in the Appendix Section A.3. The DDQN model was selected due to its superior performance over the PPO model - results are presented in Table 3.
Evaluation Scenarios and Baselines
We evaluate our framework’s performance and scalability across three distinct grid configurations, representing two primary map types: maze and warehouse. All evaluations use ten randomly generated problem instances per configuration. For the maze map type, we test on an $11\times 11$ grid with approximately 45% static obstacles (5 to 20 agents) and a larger $21\times 21$ grid with approximately 35% static obstacles (32, 64, and 96 agents). For the warehouse map type, we utilize a $25\times 25$ grid layout with approximately 24% static obstacles as shelves (32, 64, and 96 agents). Our proposed hybrid framework is evaluated through two variants, Alert-BFS and Alert-A*, which are named based on their RL policy rollout strategy (Breadth-First Search or weighted A-star search, respectively). The impact of our specific replanning strategies is further analyzed in an ablation study in Appendix Table 5. These are compared against search-based planners, Conflict-Based Search (CBS) (Sharon et al. 2015) and Improved CBS (ICBS) (Boyarski et al. 2015), and leading learning-based online planners PRIMAL (Damani et al. 2021a) and SCRIMP (Wang et al. 2023) with the MAPF formulation of disappear at goal. While our framework variants operate within our proposed four-stage pipeline (S1–S4), the baseline methods are executed as standalone algorithms. Per-instance time limits are 50 seconds for $11\times 11$ mazes, 30 minutes for $21\times 21$ mazes, and 1 hour for $25\times 25$ warehouse scenarios. For online planners, the step limits are 256 for the $11\times 11$ maze, and 512 for both the $21\times 21$ maze and the $25\times 25$ warehouse, as detailed in Table 1. Additional results on various other map configurations are provided in the Appendix Section A.4.
@@
We evaluate our MAPF framework using standard metrics. Success Rate (SR) is the proportion of instances solved within defined time limits. For successful instances, Makespan (MS) measures the time until the last agent reaches its goal. The number of Collisions (CO) measures the total count of path disagreements detected and successfully resolved by the framework to produce the final collision-free solution. A non-zero CO in a successful run reflects the planner’s total conflict resolution workload. Finally, total Time (T) measures the computational cost until a solution is found or the time limit is exceeded. Learning-based methods are evaluated by the average number of steps taken in the environment.

### v2

@@ 3.2 Proposed Approach for IC–MAPF
where $v^{k}_{0}=s_{k}$, $v^{k}_{\tau_{k}}=g_{k}$, and $\pi_{\theta}$ is a parameterized RL policy trained to minimize path length and collision risk over a planning horizon $H$. No information about other agents is exchanged during this stage. Let $\tau_{i}$ denote the set of makespan of each agent and $\tau=\{\tau_{1},\dots,\tau_{n}\}$ be the set of all such makespans.
@@
S4: Replanning
Upon receiving a collision alert for conflict time $t_{j}$ and rewind parameter $r$, the selected agent $a_{c_{k}}$ conceptually decomposes its original trajectory $\rho_{c_{k}}=(v^{c_{k}}_{0},\dots,v^{c_{k}}_{\tau_{c_{k}}})$ into a fixed prefix, a replanning window, and an untouched suffix. We denote the fixed prefix up to just before the rewind step as
@@
All S4 operators then construct a new path segment $\rho_{c_{k}}^{\prime}$ that replaces the original states on the interval $\{t_{j-r},...,t_{j+r}\}$.
(S4.0) Yield-based local coordination. As a first step, the controller attempts a lightweight yield maneuver. If the conflict set $A_{c}$ contains an agent that is already parked at its start or has reached and is waiting at its goal at time $t_{j}$, we mark this agent as anchored and select it as the replanning agent $a_{c_{k}}$; otherwise, S3 selects the replanning agent $a_{c_{k}}\in A_{c}$ using a heuristic policy (e.g., farthest from goal). In both cases, the remaining agents in $A_{c}$ are treated as protected and keep their current plans. Starting from its rewind state $v^{c_{k}}_{t_{j-r}}$, $a_{c_{k}}$ then executes a short-horizon yield: it moves to a nearby parking cell, waits while the protected agent(s) traverse the conflict region, and then returns to a state from which it can resume progress toward $g_{c_{k}}$. The parking cell is chosen by a local search that ensures the resulting detour is free of vertex and edge conflicts with all other agents over the considered window. If no such parking cell exists, or if the yield trajectory does not eliminate the conflict when re-simulated by the controller, the plan is abandoned, and the system falls back to try the next strategy.
For static obstacle avoidance (S4.1), the constraint set, $\Delta_{c}$ in Eq. 6, is fixed by the alert and encodes the static cells that must be avoided. The “bounded” aspect comes from limiting how much of the trajectory is recomputed in time: we select a sub-initial point $v^{c_{k}}_{t_{j-r}}$ and a sub-goal (either the true goal $g_{c_{k}}$ or $v^{c_{k}}_{t_{j+r}}$), and only replan this suffix under the fixed constraint. The segment of the new path is generated as:
$\rho_{c_{k}}^{\prime}=\pi_{\theta}(v^{c_{k}}_{t_{j-r}},v^{c_{k}}_{t_{j+r}}\mid v^{c_{k}}_{t_{i}}\notin\Delta_{c},\forall t\in\{t_{j-r},...,t_{j+r}\}).$ (6)
@@
If bounded static replanning still fails to resolve the conflict, the controller may share richer information and trigger dynamic obstacle avoidance (S4.2), where the replanning must account for the predicted movements of other agents involved in the collision. In this case, the RL policy is conditioned on the sub-paths of the conflicting agents as illustrated by the constraints in Eq. 7
$\rho_{c_{k}}^{\prime}=\pi_{\theta}(v^{c_{k}}_{t_{j-r}},v^{c_{k}}_{t_{j+r}}\mid v^{c_{k}}_{t}\neq v^{c_{l}}_{t},v^{c_{l}}_{t}\in\rho_{c_{l}},$
$\forall l\neq k,a_{c_{l}}\in A_{c},\forall t\in\{t_{j-r},...,t_{j+r}\}).$ (7)
@@
The information guiding the replanning–whether it constitutes minimal details such as static obstacle constraints (Eq. 6) or more detailed sub-path information about colliding agents’ paths treated as dynamic obstacles (Eq. 7)–is integrated into the RL agent’s decision-making process (by modifying its state representation, more details in Section 3.4) to guide it towards a conflict-free solution.
In all cases, the agent’s new complete trajectory $\rho^{new}_{c_{k}}$ is formed by concatenating the initial segment with the newly planned one, given by
$\rho^{new}_{c_{k}}=\rho_{c_{k}|t_{j-r-1}}\,\|\,\rho_{c_{k}}^{\prime}\,\|\,\rho_{c_{k}|t_{j+r+1}}.$ (8)
Each updated trajectory is submitted to the central controller. The iterative cycle of detection (S2), control (S3), and decentralized replanning (S4) continues until no further conflicts can be resolved by single-agent strategies or a global termination condition is reached.
In a small fraction of cases, even dynamic single-agent replanning leaves a persistent local conflict among a tightly coupled subset of agents. To address these situations, the controller invokes a local joint planner (S4.3) over a subset $A_{c}^{J}\subseteq A_{c}$ and a bounded time window around the conflict. Let $\{t_{j-r_{J}},...,t_{j+r_{J}}\}$ denote a short horizon centered at $t_{j}$. For each agent $a_{k}\in A_{c}^{J}$, we treat $v^{k}_{t_{j-r_{J}}}$ as its joint-planning start state and $v^{k}_{t_{j+r_{J}}}$ as a temporary sub-goal, and we solve a small-horizon problem in the joint configuration space of $A_{c}^{J}$ over this interval while treating all other agents as fixed reservations. The resulting joint segment $\{\rho_{k}^{\mathrm{joint}}\}_{a_{k}\in A_{c}^{J}}$ is spliced into the original trajectories between $t_{j-r_{J}}$ and $t_{j+r_{J}}$, analogous to Eq. 8.
If none of the above strategies can successfully remove a conflict, the corresponding agent is temporarily deferred: it is parked (typically at its start) and excluded from further replanning during the current phase. Once the non-deferred agents have reached a conflict-free configuration up to the current makespan, a second phase is initiated in which the deferred agents are reintroduced and planned using the same tiered strategies. This two-phase design specifically targets deadlock-like situations by decoupling stubborn local clusters and allowing them to be resolved after the rest of the system has stabilized.
3.3 Execution Flow
Our hybrid MAPF framework manages a precise flow of information, which is central to our goal of reducing inter–agent communication while preserving solution quality. The process begins with fully decentralized agent planning (S1), where each RL-driven agent uses only its local information to compute an intended path $\rho_{k}$ and submits this plan to a central coordinator. The coordinator, leveraging its global view of all submitted paths, performs collision detection (S2) to identify the set of potential conflicts $C(\rho,\tau)$. For each conflict, the control module (S3) issues a targeted alert $\mathcal{A}(c)$: it selects a replanning agent $a_{c_{k}}$, chooses a rewind window $\{t_{j-r},..,t_{j+r}\}$, and specifies which tiered strategy should be invoked.
In S4, the selected agent first attempts to resolve the conflict using minimal information through a local yield maneuver (S4.0), which relies only on occupancy checks and a nearby parking cell. If the yield is not applicable or fails, S3 gradually increases the information budget: it provides a static constraint set derived from the conflict (S4.1), then, if needed, short sub-paths of conflicting agents are treated as dynamic obstacles over the same window (S4.2). When these single-agent repairs are insufficient, a local joint A* over a small group of agents and a bounded horizon is invoked (S4.3), with all remaining agents treated as fixed reservations. Agents that still cannot be repaired are temporarily deferred and handled in a second phase with relaxed attempt limits to break deadlocks.
As shown in Appendix A.2, each tiered strategy in S4 is sound: whenever it returns a modified plan, the resulting joint execution remains collision-free over its replanned window. Moreover, under standard MAPF assumptions on finite grids with individually feasible start-goal pairs, the aggregated strategy with the deferral mechanism yields a complete repair procedure for our setting: every resolvable conflict is eventually eliminated, or else identified as belonging to an irreducible deadlock component among deferred agents. This tiered execution loop ensures that additional information is only shared on demand, allowing our framework to maintain feasibility while operating with significantly reduced information exchange.
@@ 3.4 Policy Representation
Our framework employs an RL-based decentralized planner for dynamic replanning S4.2 that incorporates collision awareness directly into the agent’s observation, enabling effective path planning in dynamic multi-agent environments.
Observation Space: Each agent observes a tensor $\mathbf{s}\in\mathbb{R}^{H\times W\times 4}$, where the channels encode: (a) a binary ObstacleMap for static obstacles, (b) an AgentMap for the agent’s current position, (c) a GoalMap denoting the goal location, and (d) an AlertMask, initially zero and updated online to reflect collision alerts generated by the centralized detection module. The observation is augmented with low-dimensional features consisting of a unit vector toward the goal and the Euclidean distance to the goal.
Action Space: The agent operates in a discrete action space $\mathcal{A}=\{0,1,2,3,4\}$, corresponding to movements in the four cardinal directions and a WAIT action.
Reward Structure: The reward function, following standard in learning-based approaches, encourages efficient, collision-free navigation. Agents receive a reward of +20 upon reaching the goal and are penalized for obstacle collisions (-3), timeouts (-2), each timestep (-0.02), and waiting actions (-0.1). An additional penalty of -0.05 is applied when the agent is near a dynamic obstacle.
@@
The collision detection module (S2) functions as a deterministic, rule-based system. It takes the set of all current agent trajectories $\rho=\{\rho_{1},\dots,\rho_{n}\}$ as input. For each timestep $t=\{0,1,...,T_{M}\}$, where $T_{M}$ is maximum makespan, the module systematically scans for vertex and edge conflicts, and reports these to the S3 control module for resolution. The detection process for all conflicts is computationally efficient, with a time complexity of $O\bigl(\sum_{k}|\rho_{k}|\bigr)$ per cycle, linear in the sum of all agent path lengths.
@@ 4 Experimental Setup
In this section, we first describe how we train the per‐agent navigation policy used in stage S1, then benchmark dataset and planner details used to evaluate stages S2–S4, and define the performance metrics used to evaluate. We additionally provide details on the hardware experiments performed.
@@
We train a parametrized Q-network $Q_{\theta}(s,a)$ using Double DQN (Van Hasselt et al. 2016) with prioritized experience replay (PER) and $\varepsilon$-greedy exploration. The policy is trained to reach a specified goal while avoiding both static and dynamic obstacles. Dynamic obstacles follow valid precomputed trajectories with hidden goals, enabling the simulation of online collision alerts during inference from the S3 stage. Training is conducted for 30,00030{,}000 episodes on an $11\times 11$ maze, with each episode capped at $T_{\max}=50$ steps. Environmental difficulty is increased through a curriculum: the first 500 episodes use static obstacle density $\rho_{s}=0.10$ with no dynamic obstacles; episodes 500–2999 use $\rho_{s}=0.10$ with one dynamic obstacle; episodes 3000–5999 use $\rho_{s}=0.20$ with two dynamic obstacles; and all remaining episodes use $\rho_{s}=0.30$ with four dynamic obstacles.
Observations are normalized to zero mean and unit variance and stored as transitions $(s_{t},a_{t},r_{t},s_{t+1})$ in a PER buffer of size $10^{6}$. Mini-batches of size 128 are sampled for learning. Action selection follows an $\varepsilon$-greedy policy, with $\varepsilon$ decayed from $\varepsilon_{0}=1.0$ to $0.01$ according to $\varepsilon_{t+1}=\max(0.01,0.999\,\varepsilon_{t})$. At all times, action selection and target computation are constrained by a validity mask $m_{t}\in\{0,1\}^{|\mathcal{A}|}$, ensuring that only feasible actions are considered (Damani et al. 2021). Network parameters are optimized using Adam with learning rate $\alpha=3\times 10^{-4}$ and discount factor $\gamma=0.97$. A separate target network with parameters $\theta^{-}$ is updated every 300 steps. The Double DQN target is computed as $y_{t}=r_{t}+\gamma\,Q_{\theta^{-}}(s_{t+1},\arg\max_{a^{\prime}}Q_{\theta}(s_{t+1},a^{\prime})\ \text{s.t.}\ m_{t}(a^{\prime})=1)$, and the temporal-difference error as $\delta_{t}=y_{t}-Q_{\theta}(s_{t},a_{t})$. Training minimizes the PER-weighted Bellman loss $L(\theta)=\mathbb{E}_{i\sim p(i)}[w_{i}\,\delta_{i}^{2}]$, where $p(i)\propto|\delta_{i}|^{\alpha}$ and $w_{i}$ are importance-sampling weights.
Simulation: Evaluation Scenarios and Baselines
We evaluate IC–MAPF on standard benchmark maps from the Moving AI Lab dataset (Stern et al. 2019b), including random grids ($32\times 32$, $64\times 64$ with 20% obstacles) and structured maps (den312d and warehouse). Agent counts range from 10–128 on random maps and 8–128 on structured maps. For each configuration, we generate 20 random instances and report aggregated results in Table 3. We compare IC–MAPF against representative search- and learning-based baselines: CBS, DCC, SCRIMP, and EPH. IC–MAPF is given a 10-minute limit per instance (Python implementation), CBS is capped at 120 seconds, and learning-based solvers are run with step limits (128 for random maps, 256 for den312d/warehouse) and a 5-minute wall-clock bound. Section 5 presents comparative results highlighting IC–MAPF’s scalability, solution quality, and communication efficiency.
@@
We evaluate our MAPF framework using standard metrics. Success Rate (SR) is the proportion of instances solved within defined time limits. For successful instances, Makespan (MS) measures the time until the last agent reaches its goal. To quantify information sharing, we define an abstract Information Unit (IU) as the data required to represent one agent’s state (e.g., its coordinates) at a single timestep. For learning-based methods, IU is estimated by counting the number of other agents observed within an agent’s field of view at each timestep, aggregated over all agents and time. This unified abstraction enables direct comparison of information load across algorithms. For IC–MAPF, each cell-level constraint issued in S3 for S4 replanning counts as one IU. For learning-based methods, IU reflects the other-agent information visible within each agent’s FOV, and we additionally count any inter-agent communication (as in DCC and EPH). We omit IU for CBS because it uses full global information.
Hardware Experiments: TurtleBot4
We evaluate our approach on five TurtleBot4 robots operating in a $6{\times}6$ indoor grid. Five problem instances are executed using a ROS 2 discovery server to ensure stable communication. Robots interface with the central controller only through a shared synchronization channel, with all other ROS topics kept isolated and no onboard perception used. At the start, each robot resets its pose via the ROS 2 reset_pos service. Execution proceeds in strict lockstep: at each step, a robot (i) performs the required in-place rotation and acknowledges completion, and then (ii) executes a fixed $0.45$ m forward motion and acknowledges again. The controller advances the system only when all robots have completed both actions, ensuring fully synchronized multi-robot execution. Details on test maps, setup images, and robot execution metrics are provided in Appendix Figures 45 and Tables 45.
Search-based Hybrid Learning-based Solvers
Map $m$ CBS IC–MAPF SCRIMP DCC EPH
[…]
[Uncaptioned image] random $\times32$-20 10-30 100 47.6 100 54.0 29.6 100 50.3 19x 100 55.5 11x 98 57.4 9.0x
40-60 65 85.3 100 98.9 161.5 95 62.1 12x 85 82.1 23x 97 77.4 22x
70-90 0 - 98 73.4 1824.2 77 83.8 2.6x 50 111.3 10x 53 114.5 12x
100 0 - 85 76.5 3553.8 60 97.2 2.0x 40 117.5 11x 50 121.0 12x
128 0 - 0 - - 70 96.6 - 0 128.0 - 5 127.5 -
[Uncaptioned image] random $\times64$-20 10-30 85 106.3 100 113.9 10.2 23 121.9 137x 97 105.9 7.9x 98 104.1 6.5x
40-60 42 130.0 100 125.6 88.5 0 128.0 - 83 116.8 9.5x 98 114.0 7.3x
70-90 0 - 98 129.1 213.5 0 128.0 - 43 125.0 14x 85 121.1 11x
100 0 - 90 130.1 462.1 0 128.0 - 25 127.2 13x 45 125.3 11x
128 0 - 5 132.0 607.6 0 128.0 - 0 128.0 - 15 127.7 20x
[…]
16 85 128.7 100 124.0 21.8 0 256.0 - 100 130.7 8.8x 100 122.2 5.9x
32 20 148.0 100 136.7 105.2 0 256.0 - 85 163.5 11x 100 127.0 6.0x
64 0 - 100 153.1 759.9 0 256.0 - 90 190.7 11x 100 146.1 8.8x
128 0 - 0 - - 0 256.0 - 5 254.8 - 90 207.7 -
[Uncaptioned image] warehouse $1×\times6$3 8 100 175.7 100 192.8 0.8 0 256.0 - 100 191.8 5.8x 100 191.6 2.3x
16 100 196.2 100 215.1 1.0 0 256.0 - 100 213.3 16x 100 212.9 11x
32 15 221.3 100 220.3 10.7 0 256.0 - 100 219.2 8.3x 100 218.1 5.4x
64 0 - 100 227.6 23.6 0 256.0 - 100 227.6 15x 100 227.8 11x
128 0 - 100 234.7 139.1 0 256.0 - 95 236.8 14x 90 239.1 11x
[…]